### v1

# MUC: Mixture of Uncalibrated Cameras for Robust 3D Human Body Reconstruction

Abstract
Multiple cameras can provide multi-view video coverage of a person. It is necessary to fuse multi-view data, e.g., for subsequent behavioral analysis, while such fusion often relies on calibration of cameras in traditional solutions. However, it is non-trivial to calibrate multiple cameras. In this work, we propose a method to reconstruct 3D human body from multiple uncalibrated camera views. First, we adopt a pre-trained human body encoder to process each individual camera view, such that human body models and parameters can be reconstructed for each view. Next, instead of simply averaging models across views, we train a network to determine the weights of individual views for their fusion, based on the parameters estimated for joints and hands of human body as well as camera positions. Further, we turn to the mesh surface of human body for dynamic fusion, such that facial expression can be seamlessly integrated into the model of human body. Our method has demonstrated superior performance in reconstructing human body upon two public datasets. More importantly, our method can flexibly support ad-hoc deployment of an arbitrary number of cameras, which has significant potential in related applications. We will release source code upon acceptance of the paper.
Keywords:
3D reconstruction Multiple cameras
footnotetext: co-first authors; corresponding author
1 Introduction
Recently, there has been a growing interest in creating 3D whole-body models from 2D images, especially for virtual reality and 3D animation. This process, known as expressive whole-body mesh recovery, combines the estimation of 3D human body pose, hand gesture, and facial expression. The progress is remarkable in single camera (monocular) setting such as SMPLer-X [3], which has become a foundation model in this field. Yet using a single camera still faces challenges such as occlusions, which can severely degrade the reconstruction quality.
Using multiple camera views helps overcome the above challenges by providing visual cues of various angles [1]. Multi-view videos, where scenes and objects are filmed simultaneously with several cameras, are common in many applications such as behavioral studies. For instance, gathering detailed multi-view whole-body data contributes to assessing children’s behaviors when playing in a naturalistic environment [28].
Given multi-view videos, it is typically the first step for many traditional methods to perform spatial calibration of cameras [7]. However, the process of multi-camera calibration is troublesome, often requiring additional calibrating toolkits and complex optimization. It is time-consuming such that multiple cameras cannot be efficiently and conveniently deployed to many sites. To this end, we will propose a new method to fuse multiple uncalibrated cameras for 3D human body reconstruction in this paper.
Most existing multi-view methods are built on single-view approaches, e.g., by directly averaging the prediction results from individual views to estimate 3D human body pose and shape [11, 12, 10]. However, directly averaging results from different views often introduces incoherence and reduces accuracy. Some other work chooses to map multiple views onto a common space to achieve feature fusion [30], which requires designing relatively complex coordinate transforms and thus increases the complexity of the method. Moreover, since multi-view data is relatively scarce compared to single-view data, it is harder to train a multi-view model directly to achieve outstanding performance.
In this study, we propose Mixture of Uncalibrated Cameras (MUC), a novel approach for reconstructing the 3D pose, shape, and facial expression of the human body from multiple uncalibrated camera views. This method is structured as a pipeline of single-view-encoding first and multi-view-fusion next. We observe that early single-view methods can learn well the inherent connection from a certain observing view to the human pose. Thus we can apply an established single-view model to each camera independently, reconstructing body parts from individual perspectives. We further observe that the body parts, if captured from a closer camera and without occlusion, often lead to more accurate reconstruction. Therefore, we propose to assess the quality of single-view reconstructions and assign dynamic weights to key body parts across different views. In this way, we can optimally fuse the multi-view data.
It is critical for our method to assess the quality of the single-view reconstuctions. To this end, we propose two strategies. (1) Joint Reweighting. For all body joints predicted from each view, we use a joint reweighting network to acquire the weights of the human body joints based on the predicted camera positions. We also implement a distance-based ranking loss function, aiming to reduce the weights of body parts that have high uncertainty in reconstruction. (2) Surface Reweighting. To integrate facial expression with body smoothly, we design a surface reweighting network. This network utilizes UV map, corresponding to vertex normals of the model of human body, to estimate dynamic weighting toward seamless combination of facial expressions and body shape.
Contributions: The contributions of this work can be summarized in three items.
  • (1)
    We present a groundbreaking multi-view method for reconstructing human pose, shape and facial expression. Our method does not require calibration or geometric mapping among views. And it is scalable to an arbitrary number of cameras (including single view) out of the box.
  • (2)
    We develop a novel ranking strategy based on distance to assess the joint importance from various viewpoints, coupled with surface reweighting for smooth and continuous fusion of facial expression and human body.
  • (3)
    Our method sets a new benchmark in calibration-free 3D human body reconstruction, outperforming existing state-of-the-art (SOTA) approaches in the field.
2 Related Work
As the demand for 3D human body reconstruction continues to increase, many efforts are focusing this field. These efforts can be categorized into single-view and multi-view methods. Benefiting from an abundance of annotated data, single-view reconstruction methods have attained notable accuracy. However, they encounter challenges when addressing human body occlusions. In contrast, multi-view methods can more adeptly handle these challenges. Yet, owing to the limited availability of multi-view datasets (i.e., datasets with multiple observation angles), considerable research is ongoing to develop effective methods for feature fusion within a multi-view framework.
Single-View Human Body Reconstruction. In the context of 3D human body reconstruction from a single viewpoint, models such as SMPL [19] and SMPL-X [24] serve as excellent low-dimensional parameterization for representing human body poses and shapes. Many existing single-view methods [6, 5, 27, 33] use these representations to learn and predict the pose and shape of human body. However, facial and finger parts are often constrained due to low resolution or small field-of-view of the image, requiring the design of independent deep learning modules to improve accuracy in these areas. For instance, ExPose [5] introduces a body-driven attention mechanism for higher-resolution crops of face and hand estimation, a dedicated refinement module, and part-specific knowledge from existing hand-only and face-only datasets. In contrast, OSX [17] proposes an end-to-end unified model to achieve pose and shape estimation for all parts of the human body, yielding outstanding results. Large quantities of high-quality real or generated human body datasets [18, 23, 8, 9], as well as non-full-body datasets [17], have played a crucial role in addressing the geometrically ill-posed problem of estimating occluded parts of the human body in a single-view scenario. SMPLer-X [3], utilizing 4.5 million images covering 32 different datasets, establishes a foundation model for human body pose and shape estimation in a single-view scenario and currently stands as SOTA on multiple test datasets.
Multi-View Human Body Reconstruction For multi-view human body reconstruction task, the first method is utilizing calibration parameters (intrinsic and extrinsic). For instance, SMPLify-X [24] extends to 3D body geometry reconstruction in a unified system. Its effectiveness is measured by multi-view image projections, focusing on minimizing 2D keypoint and silhouette reprojection errors [15]. Multi-view geometry merges features from different views to reason about self-occlusion. These visual features or predictions are transferable across views using epipolar lines, inverse projection, or a view-invariant latent space. The second way is applied when camera parameters are unavailable. A straightforward technique involves averaging the output results from a single-view model to derive the result of multi-view fusion. Additionally, several learning-based methods have been developed for this purpose. For example, a fusion module can be learned to warp a feature map from one view to align with that of another view [25], which does not rely on calibration but requires learning a large number of parameters to model cross-view geometry. Alternatively, multi-view images can be fused without calibration using the observed human body as a semantic calibration target [30]. The features are transported to the canonical body surface, where a self-attention module is used to combine multi-view features.
Figure 1: Diagram of Mixture of Uncalibrated Cameras (MUC), with the architecture of the single-view encoder shown on the top. On the right, a notation table lists the symbols and their definitions used in the model of human body.
3 Method
Our work introduces an innovative approach to reconstruct 3D human body using multi-view images from multiple uncalibrated cameras as shown in Figure 1. We adopt a pre-trained single-view Vision Transformer (ViT) based foundation model SMPLer-X [3] as encoder for each view. The core of our approach comprises two key reweighting networks: a Joint Reweighting Network (JRN) and a Surface Reweighting Network (SRN). They cleverly figures out which parts of the images are reliable and use this information to reconstruct better joints and surface.
The sequence of reweighting first the joints and then the surface is intentional, as the surface mesh is contingent upon the underlying joint-based skeleton structure. Detailed discussions of these reweighting strategies are provided in Section 3.2 for joint importance reweighting and Section 3.3 for surface importance reweighting. Upon the completion of the reweighting processes, we normalize and amalgamate all reweighted parameters into a unified parameter set. The parameter set is then fed into a decoder, which reconstructs the 3D human body.
3.1 Single-view Encoder
To exploit the benefits of extensive datasets in single-view 3D human body reconstruction, we adopt a pre-trained encoder from SMPLer-X, a generalist foundation model specialized for monocular tasks. The encoder is portrayed in the top of Figure 1.
We introduce how the single-view encoder works briefly, which is important to subsequent multi-view fusion. First, a human body image is split into a sequence of fixed-size non-overlapping patches, which are linearly projected to image tokens $T_{img}$. They are then concatenated with learnable task tokens $T_{task}$, which are transformed by ViT to $T^{\prime}_{img}$ and $T^{\prime}_{task}$. A BboxNet module predicts bounding boxes to localize face and hands in the feature map. Subsequently, MLPs are employed to regress a set of parameters $\hat{\mathcal{P}}=\{\hat{\mathbf{P}}_{body},\hat{\mathbf{P}}_{hand},\hat{\mathbf{P}}_{shape},\hat{\mathbf{P}}_{face},\hat{\mathbf{P}}_{camera}\}$. Notably, $\hat{\mathbf{P}}_{body},\hat{\mathbf{P}}_{hand}$ are joints parameter in rotation $\theta$, $\hat{\mathbf{P}}_{shape},\hat{\mathbf{P}}_{face}$ are surface parameter in PCA top-10 components. This encoder is tailored to minimize the difference between the estimated parameters $\hat{\mathcal{P}}$ and the ground-truth $\mathcal{P}$, focusing particularly on areas rich in details such as the hands and face. Further, a SMPL-X layer, as described in [24], can be the decoder to derive the 3D mesh of human body from $\hat{\mathcal{P}}$.
3.2 Multi-view Joint Reweighting
In the last section, we brief the single-view encoder. In this section, we introduce how the proposed Joint Reweighting Network (JRN) can mix the joint parameters from multiple uncalibrated cameras.
Network architecture The JRN employs two MLPs to mix joint landmark $\hat{\mathbf{P}}_{body,c}$ and hand landmark $\hat{\mathbf{P}}_{hand,c}$ based on camera position $\hat{\mathbf{P}}_{camera,c}$ for the $c$-th camera.
[…]
    $\hat{\mathbf{P}}_{body}=\frac{\sum_{c=1}^{N}\hat{\mathbf{P}}_{body,c}\cdot s_{c}^{body}}{\sum_{c=1}^{N}s_{c}^{body}}.$ (1)
[…]
    $\hat{\mathbf{P}}_{hand}=\frac{\sum_{c=1}^{N}\hat{\mathbf{P}}_{hand,c}\cdot s_{c}^{hand}}{\sum_{c=1}^{N}s_{c}^{hand}}.$ (2)
Optimization Objective Upon the previous network architecture, we propose a distance ranking loss to optimize its training. We expect that JRN can discern the relative importance of various human body joints from different camera views. Single-camera studies have shown that reconstruction errors are mostly from misjudging limbs, as more errors are associated with a larger distance from the human body to the camera. Therefore, we incorporate the distance ranking loss as a spatial piror to refine the learning process of JRN.
Figure 2: Distance ranking loss. Firstly, normalized depth maps are derived from the 3D joint predictions provided by all encoders. Then we translate these mappings into a learning-to-rank framework for optimization.
The optimization process is illustrated in Figure 2. Distance ranking loss is applied to 22 body joints, but for clarity we draw only 4 joint landmarks in Figure 2. Initially, 3D joints predicted by encoder $c$ are projected onto a 2D distance map $J_{c}\in\mathbb{R}^{21\times 3}$ in camera $c$’s image coordinate system. We then use segmentation mask, generated by RTMDet-Ins [20], to detect occluded joints. For example, in Figure 2, in camera #1’s angle, the left knee (orange dot) is occluded by right elbow. We give infinite distance values to occluded joints and normalize the distance of rest joints. In this way, JRN will be trained to neglect such occluded landmarks.
We categorize joints from different viewpoints into groups and sort them based on distance, as shown in the Figure 2, each column in the label list represents the same joint from different viewpoints. These rankings are captured as sorted label list $l\in\mathbb{N}^{N\times 21}$. Reweighting scores for these joints from $N$ different cameras ($s_{c}^{body}\in\mathbb{R}^{21}$) are collated into $s\in\mathbb{R}^{N\times 21}$, with the goal for JRN to assign higher scores to joints with top ranks across views. The learning-to-rank loss function is defined as LambdaLoss [29]:
$L_{rank}(s,l)=\sum_{l_{i}>l_{j}}\log_{2}\left(\frac{1}{1+e^{-\sigma(s_{\pi_{i}}-s_{\pi_{j}})}}\right)^{\delta_{i,j}|G_{\pi_{i}}-G_{\pi_{j}}|},$ (3)
where $\pi_{i}$ is the index of the item at position $i$ after sorting the scores, $G_{\pi_{i}}=\frac{2^{l_{\pi_{i}}}-1}{\text{maxDCG}}$ with maxDCG being the maximum discounted cumulative gain possible under given conditions, and $\delta_{i,j}=\left|\frac{1}{D_{|i-j|}}-\frac{1}{D_{|i-j|+1}}\right|$ given $D_{i}=\log_{2}(1+i)$.
Figure 3: Workflow of the Surface Reweighting Network. (a) The mixed body and hand parameters, together with the shape and expression parameters, are transformed into continuous feature maps through UV projection. (b) Employs the camera position as a condition to facilitate cross-attention operations with the feature map, resulting in the prediction of UV map-level weight maps and PCA-reduced level weight vectors.
3.3 Multi-view Surface Reweighting
In the last section, we have introduced joint reweighting network, which can mix joint landmarks from multiple cameras. This section introduces the Surface Reweighting Network (SRN), which amalgamates the body surface attached to joint from multiple cameras. However, the direct fusion of 10,475 vertices of a 3D body presents a significant challenge for current hardware capabilities and poses difficulties in optimizing the reweighting network. To address this, we propose a novel surface reweighting method that concentrates on the continuous representation of these parameters. This approach facilitates a more refined understanding of the 3D human form. The overview of SRN is illustrated in Figure 3.
As shown in Figure 3(a), initially, the mixed $\hat{\mathbf{P}}_{body}$ and $\hat{\mathbf{P}}_{hand}$ parameters, obtained from the JRN, are inputted into a SMPL-X layer alongside the $\hat{\mathbf{P}}_{shape,c}$ and $\hat{\mathbf{P}}_{face,c}$ parameters encoded from view $c$. This will generate a reconstructed human body mesh, from which vertex normals are projected onto a downsampled UV map. This projection yields the vertex normal map $VN_{shape,c}\in\mathbb{R}^{U\times V\times 3}$. Similarly, the face vertex normal map $VN_{face,c}\in\mathbb{R}^{U^{\prime}\times V^{\prime}\times 3}$ can be cropped from $VN_{shape,c}$.
We employed a conditional U-Net, which is a simplified version akin to the latent diffusion model described by Rombach et al. [26]. This U-Net is specifically informed by the predicted camera position $\hat{\mathbf{P}}_{camera,c}$. The architecture of the model is shown in Figure 3 (b). Specifically, it contains two Downsample Blocks and two Upsample Blocks with skip connections. For each block, we use $\hat{\mathbf{P}}_{camera,c}$ as K and V, and features from vertex normal map as Q to compute the cross-attention. This network predicts the weight maps $W_{shape,c}\in\mathbb{R}^{U\times V\times 3}$ and $W_{face,c}\in\mathbb{R}^{U\times V\times 3}$. These weight maps are essential for reweighting the body and facial surface features for each camera view, focusing on capturing the continuous variations in body shape and facial expressions. The final $VN_{shape}$ and $VN_{face}$ can be obtained by weighted normalization and can be supervised by the real vertex normal map generated by ground truth.
Following this, the predicted weight maps $W_{shape,c}$ and $W_{face,c}$ are converted into weight vectors $w_{shape,c}\in\mathbb{R}^{10}$ and $w_{face,c}\in\mathbb{R}^{10}$ through MLPs which bring more reasonable $\hat{\mathbf{P}}_{shape,c}$ and $\hat{\mathbf{P}}_{face,c}$ reweighting. The final $\hat{\mathbf{P}}_{shape}$ and $\hat{\mathbf{P}}_{face}$ can be obtained by weighted normalization as well.
3.4 Implementation
We train our model end-to-end by minimizing a several loss functions: $L_{smplx}$, $L_{joint2D}$, $L_{rank}$, $L_{surface}$. Each component of the loss function serves a specific purpose in the learning process. The $L_{smplx}$ loss is computed as the L1 distance between the ground truth and predicted SMPL-X parameters, providing explicit supervision for whole body joints, body shape, facial expressions, and camera position. The $L_{joint2D}$ loss is a regression loss for the projected 2D landmarks of the whole body, ensuring the accuracy of landmark localization in two-dimensional space. The $L_{rank}$ loss, as defined in Eq. 3, is used for training JRN to learn the importance of different body joints across various camera views. Finally, the $L_{surface}$ loss is calculated as the L1 distance between the ground truth and the predicted fused surface feature maps, offering direct supervision for SRN.
All model are trained using single A100 with pytorch. Adam optimizer with an initial learning rate of $1\times 10^{-5}$ for 20 epochs. The initialization weight of ViT and BboxNet from SMPLer-X [3].
Table 1: The distribution of camera numbers in the RICH dataset.
Table 2: Comparison with SMPLer-X on RICH dataset. Since SMPLer-X is a single view method, we report the mean of results from each view.
Scenes Train Validate Test
BBQ 8 N/A N/A
Gym 7 N/A 3-4
LectureHall 7 7 3-4
ParkingLot 1 8 8 N/A
ParkingLot 2 6 6 3-4
Pavallion 7 7 N/A
Method Multi-view
PA-MPJPE PA-MPVPE
SMPLer-X[3] 32.58 31.11
Ours 30.14 29.60
Table 2: Comparison with SMPLer-X on RICH dataset. Since SMPLer-X is a single view method, we report the mean of results from each view.
4 Experiments
In our experiments, we demonstrate the effectiveness and robustness of our proposed MUC method for calibration-free multi-view fusion.
  • Initially, we benchmark our method against state-of-the-art (SOTA) 3D human body reconstruction techniques, providing a detailed comparison on accuracy and robustness.
  • Subsequently, we delve into the core of our contributions, evaluating the performance with an increasing number of cameras. This involves a systematic analysis to illustrate the scalability and impact of multi-view geometry on reconstruction accuracy.
  • Lastly, we conduct an ablation study to reveal the significance of our proposed JRN and SRN based on distance ranking loss, isolating their effects and demonstrating their contributions to the overall performance.
For all experiments in this paper except ablation study, we employed the same set of model parameters trained on multi-view data to ensure our model’s adaptability to any camera count.
4.1 Datasets and Metrics
Two datasets are used to train and evaluate our method.
  • Human3.6M [9] is a large-scale multi-view dataset with ground-truth 3D human pose annotation. We follow the standard training/testing split: using subjects S1, S5, S6, S7 and S8 for training, and subjects S9 and S11 for testing. The SMPL/SMPL-X pseudo-GTs are obtained from NeuralAnnot [21]. This dataset is widely used in multiple human body reconstruction tasks and helps us to compare with other state-of-the-art methods.
  • RICH [8] is an in-the-wild and indoor multi-view dataset annotated with ground-truth 2D keypoints and 3D body mesh. We adopt the intrinsic training/testing split in the dataset. The SMPL-X pseudo-GTs are obtained from SMPLify-X. This dataset contains human body joints, shape, hand joints and facial expression annotations, which can be used to prove the effectiveness of our multi-view fusion method. The distribution of camera numbers shown in Tbale 2. For each sample in the training and validation sets, we randomly split it into two samples, each with a camera count less than or equal to 4.
Both of these datasets are based on video capture, where the positions of individuals relative to the cameras are constantly changing. In the RICH dataset, we randomly shuffle the number and combinations of cameras, hence it can be considered suitable for assessing the generalization ability of our method to camera positions.
For 3D whole-body mesh reconstruction, we utilize the mean per-vertex position error (MPVPE) and mean per-joint position error (MPJPE) as our primary metrics. In addition, we apply procrustes analysis (PA) to the recovered mesh, and report PA-MPVPE and PA-MPJPE after rigid alignment. PA-MPJPE primarily evaluates the pose estimation accuracy of the human body model, whereas PA-MPVPE evaluates the accuracy of the reconstructed 3D model. PA-MPJPE and PA-MPVPE are reported in millimeter (mm).
Table 3: Comparison between the proposed method and existing calibration-free single view and multiview methods on Human3.6M dataset. Since multiview reconstruction is not supported by single view methods, we report the mean of results from each view.
Method Camera(s) Model type Multi-view reconstruction Single-view reconstruction
PA-MPJPE PA-MPVPE PA-MPJPE PA-MPVPE
SMPLify[2] Single SMPL N/A N/A 82.3 N/A
HMR[11] Single SMPL 57.8 67.7 56.8 65.5
GraphCMR[13] Single SMPL 50.9 59.1 50.1 56.9
SPIN[12] Single SMPL 44.5 51.5 41.1 49.3
DecoMR[31] Single SMPL 42.0 50.5 39.3 47.6
Pose2Mesh[4] Single SMPL N/A N/A 47.0 N/A
I2lMeshnet[22] Single SMPL N/A N/A 41.1 N/A
PyMAF[32] Single SMPL N/A N/A 40.5 N/A
[…]
Liang[16] Multi SMPL 48.5 57.5 59.1 69.2
ProHMR[14] Multi SMPL 34.5 N/A 41.2 N/A
[…]
Ours (with translation) Multi SMPL 26.5 37.8 33.4 53.0
4.2 Comparative Analysis with SOTA Methods
In this subsection, we present a comprehensive comparison of our proposed calibration-free multi-view fusion model with current SOTA methods in the domain of 3D human body reconstruction. We employ both multi-view and single-view reconstruction methods on the Human3.6M dataset to gauge the performance of our method against an array of established techniques. The results are presented in Table 3.
Our comparison includes 3D human body reconstruction methods from 2016 to 2023. Most of them are in the category of single view, while a few are in the multi-view category. Meanwhile, many methods adopt the SMPL model, which can only represent human body pose and body shape. A more sophisticated SMPL-X, as adopted by our method, can reconstruct human body pose, hand pose, body shape and facial expression jointly. Specifically, we can treat SMPLer-X [3] as SOTA method for single-view fusion method and Yu et al. [30] as SOTA method for multi-view fusion method. To eliminate differences between using different templates, we not only evaluated using the SMPL-X template but also utilized the conversion tool provided by SMPL-X [24] to transform our prediction results into the SMPL template for evaluation.
From Table 3, we can observe the following.
  • Our Method’s Edge: Compared to other methods, our approach achieves the best multi-view reconstruction performance, with a PA-MPJPE of 31.3mm, surpassing the previous best multi-view method by 1.7mm. Additionally, it maintains a leading position in both PA-MPJPE and PA-MPVPE metrics in single-view scenarios. After converting results to the SMPL model, our method achieves even better performance, with a PA-MPJPE of 26.5mm, far ahead of any previous methods.
  • Model Bias: Under different model types (SMPL and SMPL-X), our approach demonstrates significant performance improvements in multi-view reconstruction compared to single view. Moreover, it achieves a notable enhancement in joints accuracy compared to previous methods. Since our model utilizes the SMPL-X model for both training and prediction, and the vertex count of the SMPL-X model is roughly twice that of the SMPL model, forcibly converting results to the SMPL model introduces some errors in the overall vertex alignment, thus resulting in poorer performance in the PA-MPVPE metric.
From Table 2, we can observe that our method outperforms SMPLer-X on the RICH dataset, which features more complex scenes and diverse camera settings. Our method’s visualization results are depicted in Figure 4. Compared to single-view models, our approach better compensates for the information loss caused by different viewpoints, resulting in superior reconstruction outcomes.
Figure 4: Qualitative comparison with SMPLer-X on RICH dataset. “O” stands for the original image. “S” stands for single-view reconstruction result by SMPLer-X. “F” stands for the fusion result of our method. Zoom in for better view.
4.3 Performance across Multi-view Configurations
To evaluate the effectiveness of our calibration-free multi-view fusion method, we first present a frame-based analysis for a video of three cameras. As illustrated in Figure 5, our multi-view method effectively compensates for the target’s movement in the wild, maintaining a stable PA-MPVPE fluctuation. For example, around frame #130, camera #0 exhibits incorrect pose estimation (highlighted in red bounding-box), leading to a significant spike in PA-MPVPE. Similarly, near frame #480, camera #2 shows errors in estimating arms and legs, resulting in another notable increase in PA-MPVPE for the single-view estimation. For both two cases, the fused result can correctly exhibit human body parts.
Our method is scalable to an arbitrary number of cameras out of the box, so we conduct quantitative validation to assess the impact of adding more cameras for each case. The goal is to determine how the number of cameras influences the reconstruction precision of different human body parts.
The results, depicted in Table 4, emphasize our method’s ability to leverage additional camera angles effectively. In body reconstruction, we observe a significant reduction in PA-MPJPE from 54.96mm with one camera to 30.50mm with four cameras. The PA-MPVPE also improves, decreasing from 54.58mm to 29.98mm as more cameras are incorporated. Samiliar improvement can also be observed with hand and face. Notably, the transition from one to two cameras marked the most substantial improvement, underlining the value of multi-view geometry. Although the incremental gains lessened with each additional camera, the consistent improvement across all metrics reinforces the efficacy of our approach in utilizing multiple uncalibrated views for 3D human body reconstruction in diverse settings.
Figure 5: Temporal comparison of PA-MPVPE across different camera setups over sequential frames. Mono-view reconstructions from cameras 0, 1, and 2 are depicted in varying colors, while the multi-view reconstruction is represented in red. Zoom in for better view.
Table 4: Different camera number on RICH dataset. We select cases contain 4 cameras from the test set to evaluate our method with different camera numbers.
# Cam Body Hand Face
PA-MPJPE PA-MPVPE PA-MPJPE PA-MPVPE PA-MPVPE
1 54.96 54.58 9.34 9.38 9.86
2 33.94 33.00 7.87 7.71 3.70
[…]
4 30.50 29.98 7.27 7.12 3.61
4.4 Impact of Joint and Surface Reweighting
Our qualitative evaluation focuses on the influence of the joint reweighting network on the 3D reconstruction of human body poses. Figure 6 visually demonstrates the effect of JRN’s reweighting across different camera views.
In Figure 6, the size of the red circles superimposed on the model’s joints indicates the importance level that JRN assigns to each joint. Larger circles indicate a higher score, suggesting these joints have a more significant influence on the outcome of the multi-view joint reweighting process. The first row in left, presenting a side angle view, shows only the left side of the body’s joints with minor contributions to the final reconstruction. Conversely, the second row, capturing a frontal view, displays larger weights, indicating a more substantial contribution to the final results. By emphasizing the more reliable joints and down-weighting those that are prone to error, the JRN contributes significantly to the robustness and precision of the reconstructed pose.
Figure 6: Visualization of the predicted scores by JRN across views. Larger red circle stands for larger weight. We only select 21 body joints for clear visualization. “W” stands for the visualization of joint weight. “F” stands for the fusion result of our method. Zoom in for better view.
We further perform two quantitative ablation studies experiments to objectively measure the effectiveness of the JRN and SRN: one focusing on whole body reconstruction and the other on detailed hand and face reconstruction.
Whole Body Reconstruction We utilized the Human3.6M and RICH datasets to evaluate the performance improvements. The results are compiled in Table 6. The inclusion of JRN alone, and in combination with SRN, shows a consistent decrease in PA-MPJPE and PA-MPVPE across both datasets, highlighting the benefits of the reweighting mechanisms.
Hand and Face Reconstruction The results on RICH are presented in Table 6. The quantitative analysis clearly demonstrates that the combined use of JRN and SRN yields the best performance, as indicated by the bold values for PA-MPJPE and PA-MPVPE, underscoring their importance in enhancing the precision of 3D human pose reconstruction.
Table 5: Ablation on JRN and SRN for whole body reconstruction. Best results are marked in bold.
Table 6: Ablation on JRN and SRN for hand and face on RICH dataset. Best results are marked in bold.
Human3.6M RICH
JRN SRN PA-MPJPE PA-MPVPE PA-MPJPE PA-MPVPE
33.03 36.83 34.03 30.83
[…]
Hand Face
JRN SRN PA-MPJPE PA-MPVPE PA-MPVPE
7.92 7.70 3.84
7.89 7.67 3.74
7.22 7.06 3.58
Table 6: Ablation on JRN and SRN for hand and face on RICH dataset. Best results are marked in bold.
5 Conclusion and Discussion
In this paper, we develop a new method to create 3D mesh of people from images taken by cameras that don’t need special setup and calibration. Our method is scalable to arbitrary number of cameras out of the box. This is a big step forward because it skips the complicated and time-consuming steps usually needed to get cameras ready. Our technique cleverly figures out which parts of the images are reliable and uses this information to build a better 3D reconstruction. It works by looking at how confident it is about the position of different body parts from different camera angles. By doing so, we make sure we get the best possible body pose estimation from every camera, no matter how many we use. Our method is not only easier to use in real-life situations but also gives more accurate results than current state-of-the-art methods.